Consider the following information:

  

Rate of Return if State Occurs
  State of Probability of
  Economy State of Economy Stock A Stock B Stock C
  Boom .55 .15 .22 .42
  Bust .45 .14 .04 .05

  

a.

What is the expected return on an equally weighted portfolio of these three stocks? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  Expected return %


b.

What is the variance of a portfolio invested 24 percent each in A and B and 52 percent in C? (Do not round intermediate calculations and round your answer to 6 decimal places, e.g., 32.161616.)

  

  Variance
Answer #1

1.
=0.55*(0.15+0.22+0.42)/3+0.45*(0.14+0.04-0.05)/3
=16.43333%

2.
Expected returns=0.55*(24%*0.15+24%*0.22+52%*0.42)+0.45*(24%*0.14+24%*0.04+52%*(-0.05))=0.176700

Variance=0.55*(24%*0.15+24%*0.22+52%*0.42-0.176700)^2+0.45*(24%*0.14+24%*0.04+52%*(-0.05)-0.176700)^2=0.02081475